7. A recent major study between High Blood Pressure and Heart determined the Coefficient of Non-Determination was .51   

  1. Compute the Person Product Moment Correlation Coefficient

  2. Compute the Coefficient of Non-Determination

  3. Compute the Coefficient of Non-Determination



n

1

2

3

4

5

6

7

x

11

17

13

12

15

10

12

y

21

20

20

19

18

23

20

(10 points)

X

Y

XY

X2

Y2

11

21

231

121

441

17

20

340

289

400

13

20

260

169

400

12

19

228

144

361

15

18

270

225

324

10

23

230

100

529

12

20

240

144

400

90

141

1,799

1,192

2,855

(Sum of x’s) = 90

(Sum of y’s) = 141

(Sum of x*y) = 1,799

(Sum of x2) = 1,192

(Sum of y2) = 2,855

Sample Size = 7

x = Age at First Contact with the Law

y = Age of Incarceration
Answer #1
Ʃx = 90
Ʃy = 141
Ʃxy = 1799
Ʃx² = 1192
Ʃy² = 2855
Sample size, n = 7
x̅ = Ʃx/n = 90/7 = 12.85714286
y̅ = Ʃy/n = 141/7 = 20.14285714
SSxx = Ʃx² - (Ʃx)²/n = 1192 - (90)²/7 = 34.85714286
SSyy = Ʃy² - (Ʃy)²/n = 2855 - (141)²/7 = 14.85714286
SSxy = Ʃxy - (Ʃx)(Ʃy)/n = 1799 - (90)(141)/7 = -13.8571429

Correlation coefficient, r = SSxy/√(SSxx*SSyy)

= -13.85714/√(34.85714*14.85714) = -0.6089

Coefficient of determination, r² = (SSxy)²/(SSxx*SSyy)

= (-13.85714)²/(34.85714*14.85714) = 0.3708

Coefficient of Non-Determination = 1 - 0.3708 = 0.6292